The perimeter of a badminton court is 128 feet. After a game of badminton, a player’s coach estimates that the athlete has run a total of 444 feet, which is equivalent to 6 times the court’s length plus nine times its width. What are the dimensions of the badminton court?

Answer by stanbon(48558)

The perimeter is 2l+2w=128
In addition you are told the following:
6l+9w=444
Solve by substitution:
l+w=64
l=64-w
Rewrite the 2nd equation to get:
6(64-w)+9w=444
384-6w+9w=444
3w=60
w=20 feet
l=64-20=44 feet
Cheers,
Stan H.